OpenStudy (anonymous):

write the slope intercept equation for the line with the given slope that contains the given point. show work. m=3; (6,2)

myininaya (myininaya):

y=mx+b we need to know m and b we know m=3 so y=3x+b we also know a point on this line (x=6,y=2) so 2=3(6)+b 2-18=b -16=b so y=3x-16